Back to the NIKA2 main page

Daily Reports

Bilal Ladjelate, Thomas Rose, Laurence Perotto, Stavroula Katsioli, Marco De Petris, Carsten Kramer arrive at the telescope for the 3rd NIKA2 week of the summer semester.

Oct 29th, Tuesday

16:00UT Thomas & Bilal start. Bilal gives a training session to the first-time users. One issue was noticed by AS/BL: periodic tracking glitches in Azimuth by up to +/-6". These are seen in all subscans. Pattern changes with scanning direction. Glitches are a bit reduced in height when reducing scanning speed. There is practically no wind. Elevation tracking shows no problems. However, no action was taken until Robert's mail the following morning, i.e. the pool projects were observed

21:00UT Lawrence & Stavroula take over night shift. No other issues: NIKA2, IMBfits, piic all working well. Variable weather between 0.4 and >0.8.

Oct 30th, Wednesday

05:00UT Marco & Carsten take over the morning shift. tau~0.4 (225GHz) mostly and rather stable. 091-19 continued. Pointing on nearby 0716+714. Focus on 0316+413.

around 8:00UT tau_225GHZ decreased down to 0.3. Still on the Project 091-19.

09:00UT: The periodic Az-tracking glitches are somewhat reduced in amplitude when reducing scanning speeds. For 091-19, we decide to continue scanning with 40"/s, not 60"/s for 40<El<60. A few NCS alarms: "antMD bad trace".

12:00UT: stopped. Reboot of vac1. Reset of encoders. Reboot of all mrt-lx* and vme machines became necessary.

18:00UT: Up and running again, the Az-glitches are still there. No wind. tau225~0.5 and very variable with clouds.

Oct 31st, Thursday

tau225>1. Only test observations. A possible timeline for EMIR 3mm backup projects has been prepared.

07:00: Clear sky. Only few high clouds. tau225 still above 1.0. Decision to go to EMIR. 3mm projects 051-19, 075-19, 042-19 observed with pwv~10-14mm, quite stable, mostly blue sky, some high lying ice clouds. In summary, quite OK for 3mm heterodyne work. The Az-glitches show-up in the tracking monitor. The investigation is ongoing. During the afternoon, we continue with EMIR, but during the evening, the PWV spikes and the atmosphere becomes unstable even for E0, and we switch to NIKA2 for a beammap closed window and a few tests. We stop observations once the rain arrives.

Nov 1st, Friday

06:00: NIKA2 test observations on a SNR, 1deg^2. 70"/s scanning speed. Tau high >0.7 and variable. Several thick, low clouds. Poor forecast. QL notes "sync problem" for scans 20191101s29 - s34 - all. Offline reduction works with strongly relaxed checks for NCS SYNCHRONISATION; but even with these piic warns: LOST TRACKING IN Az or SYNCHRONIZATION PROBLEM. In thick clouds. Wind increasing to ~7m/s.

08:00: tau(exp,225)=0.4. Skydip s35. Mars. Beammap with fz=0.4. Focus: fz=0.5.

09:00: tau~0.5-0.7 fluctuating. Clouds. Poor forecast. Switching to EMIR 3mm backup projects: 075-19, 042-19.

12:00: Weather degrading to 15mm pwv and more. Thick clouds. Stopping due to rain. Reboot of mrt-lx6.

Nov 2nd, Saturday

08:00 still stopped. Thick fog, rain, wind up to 18m/s. But no snow yet, temperature above zero, de-icing stays off. After input also from FP, and after having contacted JP, SS returns to the 30m to continue investigation of the az-tracking problem. One az-motor encoder exchanged. Waiting to be tested on the sky.

Nov 3rd, Sunday

09:00-14:00 still stopped. Some deicing activated during the night. Now it is switched-off again. Temperatures around zero degrees.

18:00: We had to switch to the second trap, as the 1st trap became clogged. Waiting for an hour or so for the system to stabilize again. tau225~0.25-0.4 - clear sky - no wind. The plan is to warm-up somewhat on Tuesday during maintenance, to free both traps again. The az-tracking glitches have disappeared completely, indicating that the az-motor encoder had been the culprit !

Nov 4th, Monday

03:00 tau fluctuates quite a bit: 0.25-0.5. Taumeter skydips often look poor. KID resonances move.

05:00 After some discussion and exchange of discrepant views, we decide to try observing Tony's project 091-19, despite of: tau225~0.4 rather stable, though pointing dances by 1-2". Resonances drift jointly during scans on SZ-cluster 091-19. Common atmospheric noise drifts by +/-10 to +/-30 Jy/beam @2mm, +/-40 to +/-100 Jy/beam @1mm. Blue sky. Wind <10m/s. tau rising to ~0.5.

09:00 Technical time followed by XPOL/EMIR observations till 19:00.

Nov 5th, Tuesday

Observations under poor conditions (high tau>0.5, wind near limit) till 0:40am (089-19, nikas-19), when it starts to rain and snow. Antenna stowed. De-icing is on.

DailyReportsNika2Pool1291019 (last edited 2019-11-05 08:22:04 by CarstenKramer)